11.  A gas storage tank as a volume of 4.5x10m3 at 29 °C and 101.3 kPa.  What is the new volume in cubic meters of the tank if the temperature drops to -15 °C and the pressure drops to 86 kPa? 





1
Expert's answer
2021-05-09T00:00:56-0400

Case 1:


"P=101.3kpa, V=4.5\times 10^5m^3, T=302K"


Using Ideal gas Equation-

"PV=nRT
\\[9pt]
101.3\times 10^5\times 4.5\times 10^5=nR\times 302\\[9pt]



\Rightarrow nR=1.509\times 10^{10} ~~~~~~=(1)"


Case 2:


"P=86kpa,T=258K,"


Using Ideal gas equation again-


"PV=nRT
\\[9pt]
86\times 10^5\times V=1.509\times 10^{10}\times 258\\[9pt]V=\dfrac{389.43}{86}\times 10^5

 \\[9pt]=4.528\times 10^5 m^3"